# Fourier Circuits in Neural Networks: Unlocking the Potential of Large Language Models in Mathematical Reasoning and Modular Arithmetic

### v2

@@
Fourier Circuits in Neural Networks: Unlocking the Potential of Large Language Models in Mathematical Reasoning and Modular Arithmetic
Jiuxiang Gu jigu@adobe.com. Adobe Research.    Chenyang Li lchenyang550@gmail.com. Fuzhou University.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song zsong@adobe.com. Adobe Research.    Tianyi Zhou tzhou029@usc.edu. University of Southern California.
In the evolving landscape of machine learning, a pivotal challenge lies in deciphering the internal representations harnessed by neural networks and Transformers. Building on recent progress toward comprehending how networks execute distinct target functions, our study embarks on an exploration of the underlying reasons behind networks adopting specific computational strategies. We direct our focus to the complex algebraic learning task of modular addition involving $k$ inputs. Our research presents a thorough analytical characterization of the features learned by stylized one-hidden layer neural networks and one-layer Transformers in addressing this task. A cornerstone of our theoretical framework is the elucidation of how the principle of margin maximization shapes the features adopted by one-hidden layer neural networks. Let $p$ denote the modulus, $D_{p}$ denote the dataset of modular arithmetic with $k$ inputs and $m$ denote the network width. We demonstrate that a neuron count of $m\geq 2^{2k-2}\cdot(p-1)$, these networks attain a maximum $L_{2,k+1}$-margin on the dataset $D_{p}$. Furthermore, we establish that each hidden-layer neuron aligns with a specific Fourier spectrum, integral to solving modular addition problems. By correlating our findings with the empirical observations of similar studies, we contribute to a deeper comprehension of the intrinsic computational mechanisms of neural networks. Furthermore, we observe similar computational mechanisms in the attention matrix of the one-layer Transformer. This research stands as a significant stride in unraveling their operation complexities, particularly in the realm of complex algebraic tasks.
@@ 1 Introduction
The field of artificial intelligence has experienced a significant transformation with the development of large language models (LLMs), particularly through the introduction of the Transformer architecture [108]. This advancement has revolutionized approaches to challenging tasks in natural language processing, notably in machine translation [78, 29] and text generation [60]. Consequently, models e.g., BERT [21], PaLM [18], Mistral [47], Llama [101], Llama2 [104], Llama3 [1], Gemini [98], Gemma [103], Claude3 [2], ChatGPT [75], GPT4 [76] and so on, have become predominant in NLP.
Central to this study is the question of how these advanced models transcend mere pattern recognition to engage in what appears to be logical reasoning and problem-solving. This inquiry is not purely academic; it probes the core of “understanding” in artificial intelligence. While LLMs, such as Claude3 and GPT4, demonstrate remarkable proficiency in human-like text generation, their capability to comprehend and process mathematical logic is a topic of considerable debate.
This line of investigation is crucial, given AI’s potential to extend beyond text generation into deeper comprehension of complex subjects. Mathematics, often seen as the universal language, presents a uniquely challenging domain for these models [116]. Our research aims to determine whether Transformers with attention, noted for their NLP efficiency, can also demonstrate an intrinsic understanding of mathematical operations and reasoning.
@@
This approach offers insights into why certain representations and solutions emerge from neural network training. By integrating these insights with our empirical findings, we aim to provide a comprehensive understanding of neural networks’ learning mechanisms, especially in solving the modular addition problem. We also determine the necessary number of neurons for the network to learn this Fourier method for modular addition.
Our paper’s contributions are summarized as follows:
@@
  • Neuron Count in One-Hidden-Layer Networks: We propose that in a general case, a one-hidden-layer network having $m\geq 2^{2k-2}\cdot(p-1)$ neurons can achieve the maximum $L_{2,k+1}$-margin solution. This ensures the network’s capability to effectively solve the modular addition problem in a Fourier-based method (Theorem 4.1).
  • Empirical Validation of Theoretical Findings: We validate our theoretical finding that when $m\geq 2^{2k-2}\cdot(p-1)$, for each spectrum $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$, there exists a hidden-neuron utilizes this spectrum, strong empirical support of our analysis. (Figure 1 and Figure 2).
@@ 2 Related Work
Algebraic Tasks Learning Mechanism Interpretability. The study of neural networks trained on algebraic tasks has been pivotal in shedding light on their training dynamics and inductive biases. Notable contributions include the work of [77, 35, 80] on modular addition and subsequent follow-up studies, investigations into learning parities [23, 6, 93, 95, 94, 83, 118, 112], and research into algorithmic reasoning capabilities [84, 37, 53, 63, 22, 14, 96, 71, 117, 100, 40]. The field of mechanistic interpretability, focusing on the analysis of internal representations in neural networks, has also seen significant advancements through the works of [15, 74, 68, 81, 107] and others.
@@
Theoretical Work About Fourier Transform.
To calculate Fourier transform there are two main methodologies: one uses carefully chosen samples through hashing functions (referenced in works like [34, 38, 39, 43, 42, 50, 51]) to achieve sublinear sample complexity and running time, while the other uses random samples (as discussed in [20, 82, 10, 41, 72]) with sublinear sample complexity but nearly linear running time. There are many other works studying Fourier transform [79, 66, 88, 45, 36, 61, 17, 90, 16, 91, 19, 97, 31].
@@ 3.1 Data and Network Setup
$f(\theta,x):=\sum_{i=1}^{m}\phi(\theta_{i},x),\quad\quad\phi(\theta_{i},x):=(u_{i,1}^{\top}x_{1}+\dots+u_{i,k}^{\top}x_{k})^{k}w_{i},$ (2)
where $\theta:=\{\theta_{1},\ldots,\theta_{m}\}$,  $\phi(\theta_{i},x)$ is one neuron, and $\theta_{i}:=\{u_{i,1},\dots,u_{i,k},w_{i}\}$ are the parameters of the neuron, with $u_{i,1},\dots,u_{i,k},w_{i}\in\mathbb{R}^{p}$. We use polynomial activation functions as the homogeneous requirement in Lemma 3.7 and easy sum-to-product identities calculation in Fourier analysis. Using the notation $a$ instead of the one-hot encodings $x$, we can also write:
$f(\theta,a):=\sum_{i=1}^{m}\phi(\theta_{i},a),\quad\quad\phi(\theta_{i},a):=(u_{i,1}(a_{1})+\dots+u_{i,k}(a_{k}))^{k}w_{i},$
where with $u_{i,j}(a_{j})$ being the $a_{j}$-th component of $u_{i,j}$. We consider the parameter set:
$\Theta:=\{\|\theta\|_{2,k+1}\leq 1\},\text{~{}~{}where~{}}\|\theta\|_{2,k+1}:=(\sum_{i=1}^{m}\|\theta_{i}\|_{2}^{k+1})^{1\over k+1},\text{~{}and~{}}\|\theta_{i}\|_{2}:=(\sum_{j=1}^{k}\|u_{i,j}\|^{2}_{2}+\|w_{i}\|_{2}^{2})^{1\over 2}.$
Here $\|\theta\|_{2,k+1}$ is the $L_{2,k+1}$ matrix norm of $\theta$ (Definition C.2), and $\|\theta_{i}\|_{2}$ is the $L_{2}$ vector norm of the concatenated vector of the parameters in $\theta_{i}$. The training objective over $\Theta$ is then as follows.
@@ 3.2 Margins of the Neural Networks
The minimum of the regularized objective is denoted as $\theta_{\lambda}\in\arg\min_{\theta\in\mathbb{R}^{U}}\mathcal{L}_{\lambda}(\theta)$. We define the normalized margin of $\theta_{\lambda}$ as $\gamma_{\lambda}:=h(\theta_{\lambda}/\|\theta_{\lambda}\|_{2,k+1})$. We define the maximum normalized margin as $\gamma^{*}:=\max_{\theta\in\Theta}h(\theta)$, where $\Theta=\{\|\theta\|_{2,k+1}\leq 1\}$.
Let $\mathcal{P}(D_{p})$ denote as a set containing any distributions over $D_{p}$. Then $\gamma^{*}$ can be rewritten as
$\gamma^{*}=$ $~{}\max_{\theta\in\Theta}h(\theta)=~{}\max_{\theta\in\Theta}\min_{(x,y)\in D_{p}}g(\theta,x,y)=~{}\max_{\theta\in\Theta}\min_{q\in\mathcal{P}(D_{p})}\underset{(x,y)\sim q}{\mathbb{E}}[g(\theta,x,y)],$ (3)
where the first step follows from Definition 3.4, the second step follows from Definition 3.3 and the last step follows from the linearity of the expectation.
Now, we introduce an important concept of a duality stationary pair $(\theta^{*},q^{*})$.
@@
$q^{*}$ $\in\underset{q\in\mathcal{P}(D_{p})}{\arg\min}\underset{(x,y)\sim q}{\mathbb{E}}[g(\theta^{*},x,y)],\quad\quad\theta^{*}\in\underset{\theta\in\Theta}{\arg\min}\underset{(x,y)\sim q^{*}}{\mathbb{E}}[g(\theta,x,y)].$ (4)
This means that $q^{*}$ is a distribution that minimizes the expected margin based on $\theta^{*}$, and simultaneously, $\theta^{*}$ is a solution that maximizes the expected margin relative to $q^{*}$. The max-min inequality [11] indicates that presenting such a duality adequately proves $\theta^{*}$ to be a maximum margin solution.
Recall that there is a “max” operation in Definition 3.2, which makes the swapping of expectation and summation infeasible, meaning that the expected network margin cannot be broken down into the expected margins of individual neurons. To tackle this problem, the class-weighted margin is proposed, whose intuition is similar to label smoothing. Let $\tau:D_{p}\rightarrow\Delta(\mathcal{Y})$ allocate weights to incorrect labels for every data point. Given $(x,y)$ in $D_{p}$ and for any $y^{\prime}\in\mathcal{Y}$, we have $\tau(x,y)[y^{\prime}]\geq 0$ and $\sum_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}\tau(x,y)[y^{\prime}]=1$. Then, we denote a proxy $g^{\prime}$ as the following to solve the issue.
@@
We have $g^{\prime}$ uses a weighted sum rather than max, so $g(\theta,x,y)\leq g^{\prime}(\theta,x,y)$. Following linearity of the expectation, we can get the expected class-weighted margin as
$~{}\underset{(x,y)}{\operatorname*{{\mathbb{E}}}}[g^{\prime}(\theta,x,y)]=~{}\sum_{i=1}^{m}\underset{(x,y)}{\operatorname*{{\mathbb{E}}}}[\phi(\theta_{i},x)[y]-\sum_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}\tau(x,y)[y^{\prime}]\phi(\theta_{i},x)[y^{\prime}]],$
@@ 3.3 Connection between Training and the Maximum Margin Solutions
Figure 1: Cosine shape of the trained embeddings (hidden layer weights) and corresponding power of Fourier spectrum. The two-layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We even split the whole datasets ($p^{k}=47^{4}$ data points) into the training and test datasets. Every row represents a random neuron from the network. The left figure shows the final trained embeddings, with red dots indicating the true weight values, and the pale blue interpolation is achieved by identifying the function that shares the same Fourier spectrum. The right figure shows their Fourier power spectrum. The results in these figures are consistent with our analysis statements in Lemma 4.2. See Figure 57 in Appendix I.2 for similar results when $k$ is 3 or 5.
@@
Therefore, we can replace comprehending the global minimizes by exploring the maximum-margin solution as a surrogate, enabling us to bypass complex analyses in non-convex optimization.
Furthermore, [64] states that under the following condition, the maximum-margin solutions and class-weighted maximum-margin ($g^{\prime}$) solutions are equivalent with each other.
@@
We have $g^{\prime}(\theta^{*},x,y)=g(\theta^{*},x,y)$ for all $(x,y)\in\operatorname{spt}(q^{*})$. It means: $\{y^{\prime}\in\mathcal{Y}\backslash\{y\}:\tau(x,y)[y^{\prime}]>0\}\subseteq\underset{y^{\prime}\in\mathcal{Y}\backslash\{y\}}{\arg\max}f(\theta^{*},x)[y^{\prime}].$
@@ 4 Main Result
We characterize the Fourier features to perform modular addition with $k$ input in the one-hidden-layer neuron network. We show that every neuron only focus on a distinct Fourier frequency. Additionally, within the network, there is at least one neuron for each frequency. When we consider the uniform class weighting, where $\mathcal{L}_{\lambda}(\theta)$ is based on $\tau(a_{1},\dots,a_{k})[c^{\prime}]:=1/(p-1)~{}~{}\forall c^{\prime}\neq a_{1}+\dots+a_{k},$ we have the following main result:
Theorem 4.1 (Main result, informal version of Theorem H.2).
@@
  • The maximum $L_{2,k+1}$-margin for a given dataset $D_{p}$ is:
@@
    $u_{i}(a_{i})$ $=\beta\cdot\cos(\theta_{u_{i}}^{*}+2\pi\zeta a_{i}/p),~{}~{}\forall i\in[k]\quad{\text{ and }}\quad w(c)=\beta\cdot\cos(\theta_{w}^{*}+2\pi\zeta c/p),$
@@
Theorem 4.1 tells us when the number of neurons is large enough, e.g., $m\geq 2^{2k-1}\cdot\frac{p-1}{2}$ (the lower bound of $m$ may not be the tightest in our analysis), the one hidden neural network will exactly learn all Fourier frequencies/spectrum/basis to recover the modular addition operation. More specifically, each neuron will only focus on one Fourier frequency. Thus, our analysis provides a comprehensive understanding of why neural networks trained by SGD prefer to learn Fourier-based circuits. We provide the proof sketch of Theorem 4.1 in the following section.

### v3

# Fourier Circuits in Neural Networks and Transformers: A Case Study of Modular Arithmetic with Multiple Inputs

@@
Fourier Circuits in Neural Networks and Transformers: A Case Study of Modular Arithmetic with Multiple Inputs
Chenyang Li lchenyang550@gmail.com. Fuzhou University.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Tianyi Zhou tzhou029@usc.edu. University of Southern California.
In the evolving landscape of machine learning, a pivotal challenge lies in deciphering the internal representations harnessed by neural networks and Transformers. Building on recent progress toward comprehending how networks execute distinct target functions, our study embarks on an exploration of the underlying reasons behind networks adopting specific computational strategies. We direct our focus to the complex algebraic learning task of modular addition involving $k$ inputs. Our research presents a thorough analytical characterization of the features learned by stylized one-hidden layer neural networks and one-layer Transformers in addressing this task. A cornerstone of our theoretical framework is the elucidation of how the principle of margin maximization shapes the features adopted by one-hidden layer neural networks. Let $p$ denote the modulus, $D_{p}$ denote the dataset of modular arithmetic with $k$ inputs and $m$ denote the network width. We demonstrate that a neuron count of $m\geq 2^{2k-2}\cdot(p-1)$, these networks attain a maximum $L_{2,k+1}$-margin on the dataset $D_{p}$. Furthermore, we establish that each hidden-layer neuron aligns with a specific Fourier spectrum, integral to solving modular addition problems. By correlating our findings with the empirical observations of similar studies, we contribute to a deeper comprehension of the intrinsic computational mechanisms of neural networks. Furthermore, we observe similar computational mechanisms in attention matrices of one-layer Transformers. Our work stands as a significant stride in unraveling their operation complexities, particularly in the realm of complex algebraic tasks.
@@ 1 INTRODUCTION
The field of artificial intelligence has experienced a significant transformation with the development of large language models (LLMs), particularly through the introduction of the Transformer architecture [96]. This advancement has revolutionized approaches to challenging tasks in natural language processing, notably in machine translation [70, 28] and text generation [55]. Consequently, models e.g., Mistral [40], Llama [1], Gemini [89], Gemma [93], Claude3 [2], GPT4 [68] and so on, have become predominant in NLP.
Central to this study is the question of how these advanced models transcend mere pattern recognition to engage in what appears to be logical reasoning and problem-solving. This inquiry is not purely academic; it probes the core of “understanding” in artificial intelligence. While LLMs, such as Claude3 and GPT4, demonstrate remarkable proficiency in human-like text generation, their capability to comprehend and process mathematical logic is a topic of considerable debate. This line of investigation is crucial, given AI’s potential to extend beyond text generation into deeper comprehension of complex subjects. Mathematics, often seen as the universal language, presents a uniquely challenging domain for these models [104]. Our research aims to determine whether Transformers with attention, noted for their NLP efficiency, can also demonstrate an intrinsic understanding of mathematical operations and reasoning.
@@
This approach offers insights into why certain representations and solutions emerge from neural network training. By integrating these insights with our empirical findings, we aim to provide a comprehensive understanding of neural networks’ learning mechanisms, especially in solving the modular addition problem. We also determine the necessary number of neurons for the network to learn this Fourier method for modular addition. Our paper’s contributions are summarized as follows:
@@
  • Neuron Count in One-Hidden-Layer Networks: We propose that in a general case, a one-hidden-layer network having $m\geq 2^{2k-2}\cdot(p-1)$ neurons can achieve the maximum $L_{2,k+1}$-margin solution, each hidden neuron aligning with a specific Fourier spectrum. This ensures the network’s capability to effectively solve the modular addition in a Fourier-based method (Theorem 4.1).
  • Empirical Validation of Theoretical Findings: We validate our theoretical finding that: when $m\geq 2^{2k-2}\cdot(p-1)$, for each spectrum $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$, there exists a hidden-neuron utilizes this spectrum. It strongly supports our analysis. (Figure 1 and Figure 2).
@@ 2 RELATED WORK
Algebraic Tasks Learning Mechanism Interpretability. The study of neural networks trained on algebraic tasks has been pivotal in shedding light on their training dynamics and inductive biases. Notable contributions include the work of [69, 31, 71] on modular addition and subsequent follow-up studies, investigations into learning parities [22, 6, 84, 86, 85, 73, 106, 100], and research into algorithmic reasoning capabilities [74, 33, 47, 57, 21, 14, 87, 64, 105, 91, 34]. The field of mechanistic interpretability, focusing on the analysis of internal representations in neural networks, has also seen significant advancements through the works of [15, 67, 61, 72, 95, 20, 78, 46, 17, 54] and others.
@@ 3.1 Data and Network Setup
$f(\theta,x):=$ $~{}\sum_{i=1}^{m}\phi(\theta_{i},x),$ (2)
$\phi(\theta_{i},x):=$ $~{}(u_{i,1}^{\top}x_{1}+\dots+u_{i,k}^{\top}x_{k})^{k}w_{i},$
where $\theta:=\{\theta_{1},\ldots,\theta_{m}\}\in\mathbb{R}^{(k+1)\times p}$,  $\phi(\theta_{i},x)$ is one neuron, and $\theta_{i}:=\{u_{i,1},\dots,u_{i,k},w_{i}\}$ are the parameters of the neuron with $u_{i,1},\dots,u_{i,k},w_{i}\in\mathbb{R}^{p}$. We use polynomial activation functions due to the homogeneous requirement in Lemma 3.7 and easy sum-to-product identities calculation in Fourier analysis. Using the notation $a$ instead of the one-hot encodings $x$, we can also write:
$f(\theta,a):=$ $~{}\sum_{i=1}^{m}\phi(\theta_{i},a),$
$\phi(\theta_{i},a):=$ $~{}(u_{i,1}(a_{1})+\dots+u_{i,k}(a_{k}))^{k}w_{i},$
where with $u_{i,j}(a_{j})$ being the $a_{j}$-th component of $u_{i,j}$. We consider the parameter set: $\Theta:=\{\|\theta\|_{2,k+1}\leq 1\},\text{~{}~{}where~{}}\|\theta\|_{2,k+1}:=(\sum_{i=1}^{m}\|\theta_{i}\|_{2}^{k+1})^{\frac{1}{k+1}},$ and $\|\theta_{i}\|_{2}:=(\sum_{j=1}^{k}\|u_{i,j}\|^{2}_{2}+\|w_{i}\|_{2}^{2})^{\frac{1}{2}}.$
Here $\|\theta\|_{2,k+1}$ is the $L_{2,k+1}$ matrix norm of $\theta$ (Definition D.2), and $\|\theta_{i}\|_{2}$ is the $L_{2}$ vector norm of the concatenated vector of the parameters in $\theta_{i}$. The training objective over $\Theta$ is then as follows.
@@ 3.2 Margins of the Neural Networks
The minimum of the regularized objective is denoted as $\theta_{\lambda}\in\arg\min_{\theta\in\mathbb{R}^{U}}\mathcal{L}_{\lambda}(\theta)$. We define the normalized margin of $\theta_{\lambda}$ as $\gamma_{\lambda}:=h(\theta_{\lambda}/\|\theta_{\lambda}\|_{2,k+1})$ and the maximum normalized margin as $\gamma^{*}:=\max_{\theta\in\Theta}h(\theta)$, where $\Theta=\{\|\theta\|_{2,k+1}\leq 1\}$.
Let $\mathcal{P}(D_{p})$ denote a set containing all distributions over $D_{p}$. Then $\gamma^{*}$ can be rewritten as
$\gamma^{*}=$ $~{}\max_{\theta\in\Theta}h(\theta)=~{}\max_{\theta\in\Theta}\min_{(x,y)\in D_{p}}g(\theta,x,y)$
$=$ $~{}\max_{\theta\in\Theta}\min_{q\in\mathcal{P}(D_{p})}\underset{(x,y)\sim q}{\mathbb{E}}[g(\theta,x,y)],$ (3)
where the first step is from Definition 3.4, the second step is from Definition 3.3 and the last step is from the linearity of the expectation. Now, we introduce an important concept of a duality stationary pair $(\theta^{*},q^{*})$.
@@
$q^{*}$ $\in\underset{q\in\mathcal{P}(D_{p})}{\arg\min}\underset{(x,y)\sim q}{\mathbb{E}}[g(\theta^{*},x,y)],$ (4)
$\theta^{*}$ $\in\underset{\theta\in\Theta}{\arg\min}\underset{(x,y)\sim q^{*}}{\mathbb{E}}[g(\theta,x,y)].$
This means that $q^{*}$ is a distribution that minimizes the expected margin based on $\theta^{*}$, and simultaneously, $\theta^{*}$ is a solution that maximizes the expected margin relative to $q^{*}$. The max-min inequality [11] indicates that presenting such a duality adequately proves $\theta^{*}$ to be a maximum margin solution. Recall that there is a “max” operation in Definition 3.2, which makes the swapping of expectation and summation infeasible, meaning that the expected network margin cannot be broken down into the expected margins of individual neurons. To tackle this problem, the class-weighted margin is proposed, whose intuition is similar to label smoothing. Let $\tau:D_{p}\rightarrow\Delta(\mathcal{Y})$ allocate weights to incorrect labels for every data point. Given $(x,y)$ in $D_{p}$ and for any $y^{\prime}\in\mathcal{Y}$, we have $\tau(x,y)[y^{\prime}]\geq 0$ and $\sum_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}\tau(x,y)[y^{\prime}]=1$. We denote a proxy $g^{\prime}$ as the following to solve the issue.
@@
We have $g^{\prime}$ uses a weighted sum rather than max, so $g(\theta,x,y)\leq g^{\prime}(\theta,x,y)$. Following linearity of expectation, we get the expected class-weighted margin as
$\underset{(x,y)}{\operatorname*{{\mathbb{E}}}}[g^{\prime}(\theta,x,y)]=$ $~{}\sum_{i=1}^{m}\underset{(x,y)}{\operatorname*{{\mathbb{E}}}}\Big{[}\phi(\theta_{i},x)[y]$
$~{}-\sum_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}\tau(x,y)[y^{\prime}]\phi(\theta_{i},x)[y^{\prime}]\Big{]},$
@@ 3.3 Connection between Training and the Maximum Margin Solutions
Figure 1: Cosine shape of the trained embeddings (hidden layer weights) and corresponding power of Fourier spectrum. The two-layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We even split the whole datasets ($p^{k}=47^{4}$ data points) into the training and test datasets. Every row represents a random neuron from the network. The left figure shows the final trained embeddings, with red dots indicating the true weight values, and the pale blue interpolation is achieved by identifying the function that shares the same Fourier spectrum. The right figure shows their Fourier power spectrum. The results in these figures are consistent with our analysis statements in Lemma 4.2. See Figure 57 in Appendix J.2 for similar results when $k$ is 3 or 5.
@@
Therefore, to comprehend the global minimize, we can explore the maximum-margin solution as a surrogate, enabling us to bypass complex analyses in non-convex optimization. Furthermore, [58] states that under the following condition, the maximum-margin solutions and class-weighted maximum-margin ($g^{\prime}$) solutions are equivalent with each other.
@@
We have $g^{\prime}(\theta^{*},x,y)=g(\theta^{*},x,y)$ for all $(x,y)\in\operatorname{spt}(q^{*})$, where $\operatorname{spt}$ is the support. It means: $\{y^{\prime}\in\mathcal{Y}\backslash\{y\}:\tau(x,y)[y^{\prime}]>0\}\subseteq\underset{y^{\prime}\in\mathcal{Y}\backslash\{y\}}{\arg\max}f(\theta^{*},x)[y^{\prime}].$
@@ 4 MAIN RESULT
We characterize the Fourier features to perform modular addition with $k$ input in the one-hidden-layer neuron network. We show that every neuron only focuses on a distinct Fourier frequency. Additionally, within the network, there is at least one neuron for each frequency. When we consider the uniform class weighting, where $\mathcal{L}_{\lambda}(\theta)$ is based on $\tau(a_{1},\dots,a_{k})[c^{\prime}]:=1/(p-1)~{}~{}\forall c^{\prime}\neq a_{1}+\dots+a_{k},$ we have the following main result:
Theorem 4.1 (Main result, informal version of Theorem I.2).
@@
  • The maximum $L_{2,k+1}$-margin for a dataset $D_{p}$ is:
@@
    $u_{i}(a_{i})=$ $~{}\beta\cdot\cos(\theta_{u_{i}}^{*}+2\pi\zeta a_{i}/p),~{}~{}\forall i\in[k]$
    $w(c)=$ $~{}\beta\cdot\cos(\theta_{w}^{*}+2\pi\zeta c/p),$
@@
Proof sketch of Theorem 4.1.
See formal proof in Appendix I.2. By Lemma 4.2, we get $\gamma^{*}$ and the single-neuron class-weighted maximum-margin solution set $\Omega_{q}^{{}^{\prime}*}$. By satisfying Condition 3.8, we know it is used in the maximum-margin solution. By Lemma 4.3, we can construct the network $\theta^{*}$ that uses neurons in $\Omega_{q}^{{}^{\prime}*}$. By Lemma E.2, we know that it is the maximum-margin solution. Finally, by Lemma H.2, we know that all frequencies are covered. ∎
Theorem 4.1 tells us when the number of neurons is large enough, e.g., $m\geq 2^{2k-1}\cdot\frac{p-1}{2}$ (the lower bound of $m$ may not be the tightest in our analysis), the one hidden neural network will exactly learn all Fourier spectrum/basis to recover the modular addition operation. More specifically, each neuron will only focus on one Fourier frequency. Our analysis provides a comprehensive understanding of why neural networks trained by SGD prefer to learn Fourier-based circuits.
Figure 2: All Fourier spectrum frequencies being covered and the maximum normalized power of the embeddings (hidden layer weights). The one-hidden layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We denote $\widehat{u}[i]$ as the Fourier transform of $u[i]$. Let $\max_{i}|\widehat{u}[i]|^{2}/(\sum|\widehat{u}[j]|^{2})$ be the maximum normalized power. Mapping each neuron to its maximum normalized power frequency, (a) shows the final frequency distribution of the embeddings. Similar to our construction analysis in Lemma 4.3, we have an almost uniform distribution over all frequencies. (b) shows the maximum normalized power of the neural network with random initialization. (c) shows, in frequency space, the embeddings of the final trained network are one-sparse, i.e., maximum normalized power being almost 1 for all neurons. This is consistent with our max-margin analysis results in Lemma 4.3. See Figure 6 and  8 in Appendix J.2 for results when $k$ is 3 or 5.